OpenStudy (anonymous):

How much money invested at 5% compounded continuously for 3 years will yield $820? Answers a. $952.70 b. $818.84 c. $780.01 d. $705.78 i got A... is that right?

OpenStudy (anonymous):

@hartnn am I right on this one?

hartnn (hartnn):

no...what formula did you use ?

hartnn (hartnn):

use, \( \large A = P (1+r/100)^n \) A = 820 n=3 r=5

hartnn (hartnn):

find P

OpenStudy (anonymous):

i used 

hartnn (hartnn):

no no, compound interest formula is \(\large A = P (1+r/100)^n\)

OpenStudy (anonymous):

oh okay so then i got like 3,3903.0 and thats not right

hartnn (hartnn):

?? how ? i am getting an answer which is in the choices...calculate again ?

hartnn (hartnn):

A = 820 n=3 r=5 find P

hartnn (hartnn):

1+r/100 = 1+5/100 = 1.05